A multiple choice quiz consists of five questions, each with three possible answers of which only one is correct. a student passes the quiz if three of more correct answer are obtained. what is the probability of a student passing the quiz by randomly guessing the answer to each question
Answer by stanbon(75887)

A multiple choice quiz consists of five questions, each with three possible answers of which only one is correct. a student passes the quiz if three of more correct answer are obtained. what is the probability of a student passing the quiz by randomly guessing the answer to each question
----------------------
Binomial Problem with n = 5 and p(correct) = 1/3
---
P(3<= x <=5) = 1 - binomcdf(5,1/3,2) = 1 - 0.7901 = 0.2099
====================
Cheers,
Stan H.
==================